Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Ongoing Systems Updates
- The outer vault door is getting more difficult to open. Something has gone askew. Will will notify Facilities via email.
- Martha Hruska is out of town this week.
- Resource Services has 3 more Dells to pick up – then all of the Dells will have been retrieved.
- Orders in limbo – the MSL project is especially, hopelessly on hold.
- Diana sent info regarding the move of E3000 machines to Education & AFA. The mailroom is to deliver them.
Software Projects Status
- Will is checking the switches to be full or duplex and set at the right speed. NT2 & the DLC server had network cards that were incorrectly set. Notify Bill if you notice any slow transmissions.
- Bill asks that Systems remove any superfluous items from NT2 if possible. NT2 is below 2 GIG in reserve for users. Win2K will have individual quotas for space limits, which will help deal with this problem in the future.
- Laptops – all have not been picked up yet. Geoffrey will send a reminder to Gary Cornwell.
- West closet air conditioning – the work order has been cut, no work schedule yet. The temperature is still over 90 degrees in the first floor Com closet. If we get any problem calls from that area, we may need to check the heat in the closet.
- FCLA backup tapes failed last month. Bobbie diagnosed it as a tape length problem. Bobbie replaced a tape; this should solve the problem.
- Net Agent – users want more statistics. This is a back burner issue.
- ILL – CLIO problems with installation on Veronica’s machine. It wouldn’t talk to Microenhancer. Winston was able to take care of the problem.
- OPS Tracker – Hilton should be finished with it this week.
- Elected Officials Database – Cynthia should be finished with it this week.
- Net Agent should be the next project undertaken, then Library Statistics. The Library Statistics project looks straightforward & should be easier.
- ILLIAD – no news.
- 462 reel and 162 cartridge Business tapes have been destroyed/removed/deleted from the tape library.
- Next step will be finding a migration path to reassign the rack/slots. Winston will work with Debra – software pieces are there, just not in place.
Back to the Systems Homepage
Last updated April 16, 2001
by Debra Harris